# How to Beat Nakamoto in the Race

### v1

@@ 8. Conclusion
This paper resolves the long-standing latency-security trade-off problems in Nakamoto consensus protocols. Central to our approach is the simplification of the blockchain representation, equating the entire blocktree structure to a concise tuple, and the discretization of the continuous-time block mining process into a discrete-time Markov decision process. These innovations have allowed to discover an optimal attack, bait-and-switch, whose exact performance is also evaluated.
Our analysis focuses on a baseline model in which the adversary is omniscient and can control message propagation delays up to a maximum of $\Delta$ time. However, we remark that the bait-and-switch attack generalizes naturally to scenarios with more constrained adversaries. In particular, even when the adversary has only partial control over the views of honest nodes, it can delay the propagation of honest blocks as much as its capabilities allow, and release the bait once it observes a second-best branch matches the public height. Determining the exact conditions under which this generalized attack remains optimal is an open question.
We further note that reinforcement learning offers a promising framework for discovering attack strategies in environments where the adversary has incomplete state information or reduced capabilities. Nonetheless, the exponential growth of the state space introduces significant computational hurdles. Our proposed reward function and compact state representation may provide a foundation for efficient reinforcement learning-based methods in future research.
@@
9. Acknowledgement
The authors sincerely appreciate the anonymous reviewers’ constructive suggestions, which improved this paper significantly.
@@ A.2. Proof of Lemma 2
which denotes the expected future value conditioned on the history $H_{n}$. It is easy to verify that $u(H_{1}),u(H_{2}),\dots$ is a martingale with respect to the filtration. Note that $u(H_{\infty})=E^{\pi^{*}}\left[\bar{G}|\mathscr{F}_{\infty}\right]$ is also well-defined. By Doob’s optional stopping theorem (durrett2010probability, ), for every stopping time $T$ and arbitrary policy $\pi$:
@@
Using (durrett2010probability, , Theorem 4.6.8), we have
@@ C.1.2. A-arrivals
We now consider the case of an A-arrival. By Lemmas 11 and 3, in the ahead state $[m,d,n,(l),A]$, we have $d\leq m\leq n$ and at least one of the following actions is optimal:
@@ D.2. Markov Chain
By (kroese1992difference, , Theorem 3.4), the moment generating function of $M$ is given by (1) and its pmf is given by (2).
@@ E.3. Pre-mining Lead
(166) $\times\int_{0}^{\infty}f_{1}((t+\Delta-g)a;n-i-j)he^{-ht}dtf_{G}(g)dg\Bigg{)}.$
@@
The age at time $s$ is srs-r, let $G$ denote srs-r. Since the jumper process is stationary at $s$, $G$ has the distribution with pdf (162).
@@
(187) $+\int_{\Delta}^{\infty}f_{1}(ga;j)\int_{0}^{\infty}f_{1}(ta;n-i-j)he^{-ht}dtf_{G}(g)dg\Bigg{)}$
Plug $P(A_{r,s}+A_{s,t}=n-i,\mathscr{J})$ into (171), we have proved (8).
@@ E.4. Probability Calculation
  • The target block is a jumper
  • The target block is not a jumper
@@
Starting from time $T_{1}$, the state transition is the same as in Appendix D. Thus, for $j>1$
(196) $P^{(j)}_{y,y^{\prime}}=\sum_{i=0}^{y^{\prime}-y}\frac{h}{\lambda}(\frac{a}{\lambda})^{i}P(W=y^{\prime}-y-i|L=j-1-y-i)$
For $j=1,\dots,k$
@@
(198) $P(Y_{k}=y|L_{t^{-}}=l,\mathscr{J})=\left[P^{(1)}\times\dots,P^{(k)}\right]_{l+1,y+1}$
@@
(207) $P(Y_{k}+M<k|L_{t}^{-}=l,\mathscr{J}^{C})$
(208) $=$ $\sum_{i=0}^{k-1}\sum_{y=l}^{k-1-i}P(M=i)P(Y_{k}=y|L_{t}^{-}=l,\mathscr{J}^{C}).$
@@
$P(Y_{k}+M<k,L_{t}^{-}=l,\mathscr{J}^{C})$
(209) $=$ $P(Y_{k}+M<k|L_{t}^{-}=l,\mathscr{J}^{C})P(L_{t}^{-}=l,\mathscr{J}^{C}),$
@@
$P(Y_{k}+M<k)$ $=\sum_{l=0}^{k-1}\Bigg{(}(P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J})$
(210) $+P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J}^{C})\Bigg{)}$
where $P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J})$ is given by (E.4) and $P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J}^{C})$ is given by (E.4). Thus, the probability of safety violation of a target height including the target transaction is given in (3), hence we finish the prove of Theorem 3.

### v2

@@ 8. Conclusion
This paper resolves the longstanding question of block safety in PoW Nakamoto consensus protocols by identifying an optimal attack strategy, termed “bait-and-switch,” and computing the precise probability of a safety violation after a given number of confirmations. Our analysis considers a baseline model where an omniscient adversary controls message propagation delays up to a maximum of $\Delta$. Central to our approach is the formulation of a tractable discrete-time Markov decision process that concisely captures the system state and the adversary’s permissible actions.
The bait-and-switch attack naturally extends to scenarios with more constrained adversaries. Specifically, when the adversary has partial control over the views of honest nodes, it can delay the propagation of honest blocks within its capabilities and release the bait once it observes a second-best branch matching the public height. Determining the precise conditions under which this generalized attack remains optimal is an open question.
Reinforcement learning offers a promising framework for identifying attack strategies in settings where the adversary has incomplete state information or reduced capabilities. However, the exponential growth of the state space introduces significant computational challenges. Our proposed reward function and compact state representation may provide a foundation for efficient reinforcement learning-based methods in future research.
@@
9. Acknowledgment
The authors sincerely appreciate the anonymous reviewers’ constructive suggestions, which improved this paper significantly. This work is supported in part by the National Science Foundation under Grant no. 2132700.
@@ A.2. Proof of Lemma 2
which denotes the expected future value conditioned on the history $H_{n}$. It is easy to verify that $u(H_{1}),u(H_{2}),\dots$ is a martingale with respect to the filtration. Note that $u(H_{\infty})=E^{\pi^{*}}\left[\bar{G}|\mathscr{F}_{\infty}\right]$ is also well-defined. By Doob’s optional stopping theorem (Durrett, 2010), for every stopping time $T$ and arbitrary policy $\pi$:
@@
Using (Durrett, 2010, Theorem 4.6.8), we have
@@ C.1.2. A-arrivals
We now consider the case of an A-arrival. By Lemmas 11 and 2, in the ahead state $[m,d,n,(l),A]$, we have $d\leq m\leq n$ and at least one of the following actions is optimal:
@@ D.2. Markov Chain
By (Kroese, 1992, Theorem 3.4), the moment generating function of $M$ is given by (1) and its pmf is given by (2).
@@ E.3. Pre-mining Lead
(165) $\times\int_{0}^{\infty}f_{1}((t+\Delta-g)a;n-i-j)he^{-ht}dtf_{G}(g)dg\Bigg{)}$
@@
The age at time $s$ is srs-r, let $G$ denote srs-r. Since the jumper process is stationary at $s$, $G$ has the distribution with pdf (161). We have
@@
(186) $+\int_{\Delta}^{\infty}f_{1}(ga;j)\int_{0}^{\infty}f_{1}(ta;n-i-j)he^{-ht}dtf_{G}(g)dg\Bigg{)}.$
Plugging $P(A_{r,s}+A_{s,t}=n-i,\mathscr{J})$ into (170), we have proved (8).
@@ E.4. Probability Calculation
  • The target block is a jumper;
  • The target block is not a jumper.
@@
Starting from time $T_{1}$, the state transition is the same as in Appendix D. Thus, for $j>1$,
(195) $P^{(j)}_{y,y^{\prime}}=\sum_{i=0}^{y^{\prime}-y}\frac{h}{\lambda}(\frac{a}{\lambda})^{i}P(W=y^{\prime}-y-i|L=j-1-y-i).$
For $j=1,\dots,k$,
@@
(197) $P(Y_{k}=y|L_{t^{-}}=l,\mathscr{J})=\left[P^{(1)}\times\dots,P^{(k)}\right]_{l+1,y+1}.$
@@
(206) $\begin{split}P&(Y_{k}+M<k\,|\,L_{t}^{-}=l,\mathscr{J}^{C})\\ &=\sum_{i=0}^{k-1}\sum_{y=l}^{k-1-i}P(M=i)P(Y_{k}=y\,|\,L_{t}^{-}=l,\mathscr{J}^{C}).\end{split}$
@@
(207) $\begin{split}P&(Y_{k}+M<k,L_{t}^{-}=l,\mathscr{J}^{C})\\ &=P(Y_{k}+M<k|L_{t}^{-}=l,\mathscr{J}^{C})P(L_{t}^{-}=l,\mathscr{J}^{C}),\end{split}$
@@
(208) $\begin{split}P(Y_{k}+M<k)=\sum_{l=0}^{k-1}&\Big{(}P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J})\\ &+P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J}^{C})\Big{)}\end{split}$
where $P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J})$ is given by (E.4) and $P(Y_{k}+M<k,L_{t^{-}}=l,\mathscr{J}^{C})$ is given by (207). Thus, the probability of safety violation of a target height including the target transaction is given in (3), hence we finish the prove of Theorem 3.